Sheridan Company has 495000 shares of $10 par value common stock outstanding. During the year Sheridan declared a 16% stock dividend when the market price of the stock was $35 per share. Three months later Sheridan declared a $0.50 per share cash dividend. As a result of the dividends declared during the year, retained earnings decreased by

$2772000 $ 466700 $3059100 $ 490050
Answer #1

Stock dividend

= 495,000*16% = 79,200

Stock dividend in cash = 79,200*35 = 2,772,000

Cash dividend = (495,000+79200)*0.50 = 287,100

Reduction in Retained earnings= 2,772,000+287,100

= 3,059,100